OpenStudy (anonymous):

Can someone explain to me how to find the right answer in mathematical terms?

OpenStudy (anonymous):

P = a + b + c

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

since the bottom of all the fractions are the same. your just adding the top value. 9x + 9x + 13x and then devide that value by 11

OpenStudy (anonymous):

you have a common denominator so you dont have to change it \[\frac{ 9x }{ 11y }+\frac{ 9x }{ 11y }+\frac{ 13x }{ 11y }=\frac{ 9x+9x+13x }{ 11y } = \frac{ 31 }{ 11y }\]

OpenStudy (anonymous):

hah fool